Find the focus and directrix of the parabola
y = 3x^2
please explain

Answer by lwsshak3(11628)

Find the focus and directrix of the parabola
y = 3x^2
please explain
***
This is a parabola that opens up.
vertex: (0,0)
axis of symmetry: x=0 or y-axis
Its basic form: x^2=4py
x^2=(1/3)y
4p=1/3
p=1/12
focus: (0,1/12) (p distance above vertex on the axis of symmetry)
directrix: y=-1/12 (p distance below vertex on the axis of symmetry)